OpenStudy (anonymous):

Multiply. (x^2+x+7)(x-7)

OpenStudy (anonymous):

\[(x2+x+7)(x−7)=x^3−6x^2−49\]

OpenStudy (mathstudent55):

Multiply every term of the left polynomial by every term of teh right polynomial. (x^2 + x + 7)(x - 7) = x^2(x) + x^2(-7) + x(x) + x(-7) + 7(x) + 7(-7) Do the multiplications, then combine like terms.

OpenStudy (anonymous):

Thank you!